The half life of Rn is 3.82 days. Determine the time it takes for 60% of the Rn to decay, the activity of 10 mg of Rn.
1
Expert's answer
2020-06-22T05:14:10-0400

Given,

"t_{1/2}=3.8\:days=3,30,048s,A_0=10^{-2}\:kg"

Formula for decay

"A(t)=A_0(1\/2)^{t\/t_{1\/2}}"

when, "60\%" decay occurs then "A(t)=40\%" of "A_0" ,thus

"0.4A_0=A_0(1/2)^{t/330048}\\
\implies t= - 330048\ln(2/5)/\ln(2)\approx4.36\times 10^5s"